OpenStudy (anonymous):

If Sam can do a job in 4 days that Lisa can do in 6 days and Tom can do in 2 days, how long would the job take if Sam, Lisa, and Tom worked together to complete it? I'm confused as to how to work the problem, I don't just want an answer.

OpenStudy (anonymous):

x((1/4) + (1/6) + (1/2)) = 1 x = 12/11 A bit more than one day.

OpenStudy (anonymous):

\[(\frac{1}{4}+\frac{1}{6}+\frac{1}{2})t=1\] solve for t

OpenStudy (saifoo.khan):

\[\Huge \frac{1}{4} + \frac{1}{6} + \frac{1}{2} = \frac{1}{x}\]

OpenStudy (anonymous):

saifoooooooooooooooooooooo!

OpenStudy (anonymous):

sattelite can u take a look at mine

OpenStudy (kinggeorge):

1/4 + 1/6 + 1/2 = 1/x where the three terms are the amount of work each person can do in a day.

OpenStudy (saifoo.khan):

Satelllllliteeee!

OpenStudy (anonymous):

\[\frac{11}{12}t=1\] \[t=\frac{12}{11}\] what pk said

OpenStudy (anonymous):

12/11 days

OpenStudy (anonymous):

Thanks!

OpenStudy (saifoo.khan):

\[\Huge \color\red {\frac{12}{11}}\]